### 10 Jan 2024 v2

# Plug-and-Play Transformer Modules for Test-Time Adaptation

Abstract
Parameter-efficient tuning (PET) methods such as LoRA, Adapter, and Visual Prompt Tuning (VPT) have found success in enabling adaptation to new domains by tuning small modules within a transformer model. However, the number of domains encountered during test time can be very large, and the data is usually unlabeled. Thus, adaptation to new domains is challenging; it is also impractical to generate customized tuned modules for each such domain. Toward addressing these challenges, this work introduces PLUTO: a Plug-and-pLay modUlar Test-time domain adaptatiOn strategy. We pre-train a large set of modules, each specialized for different source domains, effectively creating a “module store”. Given a target domain with few-shot unlabeled data, we introduce an unsupervised test-time adaptation (TTA) method to (1) select a sparse subset of relevant modules from this store and (2) create a weighted combination of selected modules without tuning their weights. This plug-and-play nature enables us to harness multiple most-relevant source domains in a single inference call. Comprehensive evaluations demonstrate that PLUTO uniformly outperforms alternative TTA methods and that selecting $\leq$5 modules suffice to extract most of the benefit. At a high level, our method equips pre-trained transformers with the capability to dynamically adapt to new domains, motivating a new paradigm for efficient and scalable domain adaptation.
1 Introduction
Pretrained transformers [7, 4, 31, 38, 39] have achieved significant recent success, however, these models are often quite large, with billions of parameters [4, 39]. This has posed challenges for their practical use, especially in edge devices, due to the computational and memory challenges of fine-tuning for downstream tasks. To address this, parameter-efficient tuning (PET) methods have been proposed including prefix/prompt-tuning [28, 27, 30], adapter [18], and LoRA [20]. In PET, the pretrained model remains frozen and only the domain-specific (extra) parameters designated for the target domain are updated based on the training data. PET has found notable success as it significantly reduces memory usage and facilitates modular adaptation of the model while often being competitive with fine-tuning.
While PET methods have demonstrated notable benefits across various domains and tasks [30, 51], their effectiveness in more challenging scenarios involving few-shot and unlabeled data or domain shifts remain as a major bottleneck [14, 39]. Specifically, we ask: How to leverage PET methods to adapt to a new domain at test-time [52] where we only observe few-shot unlabeled data from this new domain? Here, test-time adaptation (TTA) refers to adapting to new domains during testing, in an online manner, without the need for additional training data. This makes this setting more challenging than standard unsupervised domain adaptation (UDA) [47, 48]. TTA typically requires iteratively updating the source model parameters using an unsupervised objective that incorporates the new test sample from the target distribution. Existing test-time adaptation (TTA) techniques mostly depend on a single source model for adaptation [52, 53]. However, considering the dynamic nature of the test distribution, it should prove advantageous to adapt a set of source models collectively during test time [55, 46].
Table 1: Our work compared to the prior art fulfills key criteria for a comprehensive adaptation framework. To efficiently realize multiple sources, we employ PET modules within a single ViT.
Setting Source Free Adaptation On the fly Dynamic Target Multi Source
UDA
Source-free UDA
TTA
PLUTO
Figure 1: The overview of PLUTO. At test time, PLUTO efficiently combines the sources using appropriate weights determined by the current test distribution. Furthermore, we selectively update the LayerNorm (LN) parameters of the model that demonstrates the highest correlation with the test distribution.
In this work, rather than employing multiple distinct source models, we propose to plug source module(s) within a single vision transformer. We naturally can’t expect a single module to work for all target domains. Also, since target domains at test time are few-shot and there can be a large number of domains, it is impractical to train customized PET modules for each target from scratch. Instead, the ideal algorithm should identify the new domain and suitably adapt pre-existing PET modules to this domain. The main contribution of this work is providing a plug-and-play algorithm to address this challenge. We call our method PLUTO: Plug-and-pLay modUlar Test-time adaptatiOn.
Concretely, we consider a multi source domain adaptation setting and propose pretraining a PET module for a collection of source domains. These pretrained modules can then be utilized at test-time with minimal supervision and sample size. Importantly, our algorithm PLUTO benefits from having access to a module store containing a diverse set of source domains. PLUTO can retrieve the most relevant sources from this module store (70+ modules for ImageNet-C) and accurately blend them in a target-aware fashion. This procedure has two important benefits: (1) Sample and compute efficiency: Selecting a small subset of pretrained modules and weighting them is a much lower dimensional search space compared to tuning a module from scratch. (2) Inference efficiency: Unlike alternative multi-source TTA methods such as TENT-Ensemble [52], PLUTO is able to utilize multiple source domains with a single inference call.
Main Contributions. Overall, PLUTO leads to an effective strategy for harnessing not only a pretrained Vision Transformer (ViT) but also pretrained PET modules within this ViT to enhance few-shot unsupervised TTA performance. The main technical contributions of our work are:
  • PET-based TTA. To the best of our knowledge, this is the first work to employ PET for the test-time adaptation problem. We not only pretrain a large number of modules, we present innovative algorithms that can effectively select and blend an optimal subset of these modules to facilitate few-shot unsupervised adaptation to a new domain.
  • Algorithmic innovations. PLUTO facilitates multi-source TTA through weighted logit outputs of multiple modules and fine-tuning of LayerNorms (LN). Our method crucially captures the relationships between the logit outputs of multiple source modules and input instances in a single inference call and enables PLUTO to adaptively decide the benefit of each source domain. We also introduce sharpness-aware minimization (SAM) [10] for TTA and demonstrate how it boosts the TTA performance for PLUTO as well as other TTA/UDA methods.
  • Empirical impact and insights. Comprehensive evaluations demonstrate that PLUTO uniformly outperforms alternative TTA methods and that selecting $\leq$5 modules suffice to extract most of the benefit. We also provide a comparison of zero-shot vs few-shot adaptation which reveals PLUTO also has respectable zero-shot performance.
2 Related Works
2.1 Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) is widely applied in computer vision tasks, including image classification [48], semantic segmentation [47], object detection [19], and reinforcement learning [40]. These applications address distribution shift challenges by harmonizing source and target domain distributions. Methods utilize techniques like maximum mean discrepancy and adversarial learning [12, 48] for alignment. There is a growing interest in adaptation strategies that rely solely on pretrained source models due to privacy and memory storage concerns regarding source data.
2.2 Test Time Adaptation (TTA)
UDA demands substantial target data for offline adaptation, while TTA adapts batch by batch. Early work [29] used incoming test-batch stats for adaptation instead of saved training stats. TENT [52] minimizes entropy and updates batch-norm in a pretrained source model with incoming target data. DUA [33] continuously updates batch-norm stats with test batches for target alignment. TTA methods extend to segmentation tasks [49, 44, 21]. However, they often forget source knowledge. CoTTA [53] restores source knowledge stochastically to avoid excessive deviation, and EATA [34] preserves vital model weights with regularization. This paper shows that TTA ensembles mitigate forgetting.
2.3 Ensemble learning
Ensemble learning, a well-established approach, aims to create a robust and generalizable model [16]. Common ensemble techniques include voting[16], bagging[3], boosting[42, 11], and stacking[54]. These methods have found applications in various tasks, including model debiasing[9, 45], and domain adaptation[25]. Apart from considering global domain-to-domain relationships, [41, 15] introduced the use of similarity measures between examples and domains to choose data from various sources for a particular target domain. These selected data are then merged to train a single classifier. Some closely related works to ours are [25, 37], which also incorporate example-to-domain relationships but utilize an attention mechanism. They train the attention module using limited labeled data from the target domain in a supervised manner. In contrast, our method operates in an unsupervised setting and does not rely on such labeled data.
3 Problem formulation
We are given a large pretrained transformer model $f$ with size $|f|$ and input $x$. Our goal is to develop a domain adaptation technique with "module store", a set of pretrained modules $\{\theta_{j}\}_{j=1}^{N}$ of size $|\theta_{j}|\ll|f|$. Here are the steps building the store $\{\theta_{j}\}_{j=1}^{N}$:
  • Pretrain on source domains: Given a loss function $\mathcal{L}$ and a fixed set of source domains $\{\mathcal{T}_{j}\}_{j=1}^{N}$, for each $\mathcal{T}\in\{\mathcal{T}_{j}\}_{j=1}^{N}$, instead of fully fine-tuning ($f\rightarrow f_{\mathcal{T}}$) the pretrained model to the domain $\mathcal{T}$ by solving $f_{\mathcal{T}}=\text{argmin}_{f}\mathcal{L}(f;\mathcal{T})$, given the pretrained model (with modules) $f_{\theta}$, we freeze the model and only update the module $\theta=\text{argmin}_{\theta}\mathcal{L}(f_{\theta};\mathcal{T})$.
  • Adapt to a new domain: After obtaining a set of $N$ modules $\{\theta_{j}\}_{j=1}^{N}$ corresponding to $\{\mathcal{T}_{j}\}_{j=1}^{N}$, our objective is to fit a new target domain $\mathcal{T}_{new}$ without fine-tuning a new module but only utilizing the current modules. We aim to find a combination $f$ of the modules $\mathcal{G}(\{\theta_{i}\}_{i=1}^{N})$ that minimizes the loss on the new domain: $\mathcal{G}=\text{argmin}_{\mathcal{G}}\mathcal{L}(f(\mathcal{G}(\{\theta_{j}\}_{j=1}^{N}));\mathcal{T}_{new})$ A very straightfoward version of $\mathcal{G}$ is $\mathcal{G}=\sum_{j=1}^{N}w_{j}\theta_{j}$. An underlying assumption here is that the module output for new target domain $\mathcal{T}_{new}$ can be approximately expressed by the linear combination of current modules outputs.
4 Method
PLUTO’s framework involves four main steps to achieve unlabeled data adaptation:
  • We freeze the weight of the pretrained transformer, and finetune the modules on different source domains in a supervised manner separately.
  • We freeze the weight of the obtained module while keep the pretrained transformer frozen, and train the module selector in a supervised manner.
  • For a test batch at time stamp $t$, we have both the pretrained modules and pretrained models frozen while updating module selector $\mathcal{G}$ with pseudo label entropy minimization. This process helps us learn the module output combination weights and select most related source modules specific to the current batch.
  • At each time stamp $t$, we determine which source module need to update the LN affine parameters by the largest assigned weight from the updated $\mathcal{G}$.
In the second and third steps, we only require a small number of test samples to identify the source modules most relevant to the current test batch. By selecting and using these modules, we achieve results that approximate the effect of using all weighted source module outputs for forward inference. We provided an overview of PLUTO in Alg. 1
4.1 Obtain pretrained modules
Given the correspondence between $N$ source domains $S_{1},\dots,S_{N}$ and their labeled training data, by running parameter-efficient tuning on each domain separately, a set of pretrained modules can be obtained: $\{\theta_{j}\}_{j=1}^{N}$. Given a certain pretrained transformer model $f$, we denote the models equipped with pretrained source modules as $\{f_{\theta_{j}}^{j}\}_{j=1}^{N}$.
For image classification tasks, we assume a shared label space between source and target domains.
Algorithm 1 PLUTO: Plug-and-pLay modUlar Test-time adaptatiOn
1:Input: models with pretrained source modules $\{f_{\theta_{j}}^{j}\}_{j=1}^{N}$, number of modules to be selected $M$ ($M<N$), streaming sequential unlabeled test data $\{x_{i}^{(1)}\}_{i=1}^{B}\rightarrow\{x_{i}^{(2)}\}_{i=1}^{B}\rightarrow\dots\{x_{i}^{(t)}\}_{i=1}^{B}\rightarrow\dots$, Attention module $\mathcal{G}$
2:Output: $M$ scaled weights, $M$ adapted modules
3:while $t\geq 1$ do
4:     for Each $x_{i}$ in the $t$-th batch do
5:         Pass $x_{i}$ through $\{f_{\theta_{j}}^{j}\}_{j=1}^{N}$
6:         Obtain pre-softmax logits $\{l(x_{i})^{j}\}_{j=1}^{N}$
7:         Obtain weights
8:      $\{w(x_{i})^{j}\}_{j=1}^{N}=\mathcal{G}(x_{i},\{l(x_{i})^{j}\}_{j=1}^{N})$
9:         Assign pseudo-label (PL) to $x_{i}$
10:     end for
11:     Update $\mathcal{G}\rightarrow\mathcal{G^{*}}$ with PL entropy minimization on $t$-th batch (Eqn. 2)
12:     Obtain updated weights for all $x_{i}$ in $t$-th batch
13:      $\forall x_{i},\{w^{*}(x_{i})^{j}\}_{j=1}^{N}=\mathcal{G^{*}}(x_{i},\{l(x_{i})^{j}\}_{j=1}^{N})$
14:     Calculate average weights for $t$-th batch
15:      $\{\overline{w}(t)^{j}\}_{j=1}^{N}=\{\overline{\{w^{*}(x_{i})^{j}\}_{i=1}^{B}}\}_{j=1}^{N}$
16:     Find indices of top $M$ in $\{w^{*}(t)^{j}\}_{j=1}^{N}$
17:     LN update of source model with largest $\overline{w^{*}(x_{i})^{j}}$ by minimizing Eqn. 4 with Eqn. 8
18:end while
4.2 Initialization and training of module selector
Drawing from the success of the attention mechanism in achieving impressive few-shot performance in supervised model ensemble scenarios [37, 25], we have extended this observation to test-time adaptation. Even in the unsupervised context of test-time adaptation, we have found that an attention module, which acts as a module selector, can effectively integrate information from multiple source domains into the target domain by dynamically assigning weights and updating on-the-fly. This enables PLUTO to plug modules from the module store and adapt them at test time with few-shot unlabeled data.
In order to convert 2D image input into sequential data for transformers and module selector to process, the image input $x\in\mathbb{R}^{H\times W\times C}$ is reshaped into a sequence of flattened 2D patches [8]$\mathbf{X}_{p}\in\mathbb{R}^{l\times(P^{2}\cdot C)}$, where $(H,W)$ is the height and width of the original image, $C$ is the number of channels, $(P,P)$ is the size of each segmented image patch, and $l=HW/P^{2}$ is number of patches and the effective input sequence length. Assuming that the transformer uses constant latent vector size $d$ through all of its layers, we flatten the sequence of the patches and map them to $d$ dimentions: $\hat{\mathbf{X}}=[\mathbf{X}_{p}\mathbf{E}]_{p=1}^{N},\mathbf{E}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$. The patch embedding projection $\mathbf{E}$ is a (frozen) trainable parameter in the pretrained ViT.
The module selector, which is an attention module, gauges the relevance between the output logits of the source modules and instances $\bm{x}_{i}$ in the current ($t$-th) target (abbreviated as $T$) batch $D_{T}^{(t)}=\{x_{i}^{(t)}\}_{i=1}^{B}\sim\mathcal{D}_{T}^{(t)}$. It then assigns weights based on this relevance. Logits of source models, on the other hand, are a function of both the pretrained module and the pretrained transformer and are also capable of modeling the uncertainty in predictions from different sources. We perform max-pooling over the sequence $\hat{\mathbf{X}}=[\mathbf{X}_{p}\mathbf{E}]_{p=1}^{N}\in\mathbb{R}^{l\times d}$ and obtain $\hat{x}\in\mathbb{R}^{d}$ as the representation of the original image input. Furthermore, we pass the input instance $x$ through $N$ pretrained models and obtain pre-softmax logits $\{l(x_{i})^{j}\}_{j=1}^{N}\in\mathbb{R}^{v}$. With four trainable parameters: $W_{d,x}\in\mathbb{R}^{d\times d^{\prime}_{x}},W_{u,x}\in\mathbb{R}^{d^{\prime}_{x}\times d^{\prime}},W_{d,l}\in\mathbb{R}^{v\times d^{\prime}_{l}},W_{u,l}\in\mathbb{R}^{d^{\prime}_{l}\times d^{\prime}}$, the module selector first project the pre-softmax logits and the representation of the original image input into another representational space: $\mathbf{h}_{x}=W_{u,x}^{\intercal}\cdot\gamma(W_{d,x}^{\intercal}\cdot\hat{x}),\mathbf{h}_{l,j}=W_{u,l}^{\intercal}\cdot\gamma(W_{d,l}^{\intercal}\cdot{l(x)^{j}})$. With application of LN [2] we are able to obtain the final projected representations $\mathbf{h}_{x},\mathbf{h}_{l,j}\in\mathbf{R}^{d^{\prime}}$ and compute the attention weight $w(x)^{j}$ to capture the relationship between the input $x$ and its representation from the pretrained models $l(x)^{j}$:
$w(x)^{j}=\mathcal{G}(x_{i},\{l(x_{i})^{j}\}_{j=1}^{N})=\frac{e^{\mathbf{h}_{l,j}\cdot\mathbf{h}_{x}}}{\sum_{k=1}^{N}e^{\mathbf{h}_{l,k}\cdot\mathbf{h}_{x}}}$ (1)
The ensembled logit is the linear combination of the weighted logits: $l(x)=\sum_{j=1}^{N}w(x)^{j}l(x)^{j}.$ If the ground-truth (GT) label of the target sample $x$ is known (e.g. for source datasets $\{S_{j}\}_{j=1}^{N}$), the module selector $\mathcal{G}$ can be randomly initialized and trained/updated by minimizing its cross-entropy loss between $\text{softmax}(l(x))$ and the GT label $y$. As disucssed in [37], this enables $\mathcal{G}$ to capture the sample-specific preference of different source models.
Otherwise, if the GT label of the target sample $x$ is unknown, we denote its pseudo-label, as predicted by source $j$, as $\hat{y}_{j}=f^{j}_{\theta_{j}}(x)$. We linearly combine these pseudo-labels by attention weights $\{w(x)^{j}\}_{j=1}^{N}$ to get the weighted pseudo-label $\hat{y}=\sum_{j=1}^{N}w(x)^{j}\hat{y}_{j}$. Using these weighted pseudo-labels for all the samples in the current ($t$-th) batch we calculate the expected Shannon entropy as
$\mathcal{L}=-\mathbf{E}_{\mathcal{D}_{T}^{(t)}}\sum_{c=1}^{K}\hat{y}_{c}\log(\hat{y}_{c}),$ (2)
where $K$ is the number of classes and $\hat{y}_{c}$ is the $c$-th entry of the predicted pseudo-label. By performing pseudo-label entropy minimization on Eqn. 2, we can update the module selector parameters(from $\mathcal{G}$ to $\mathcal{G}^{*}$) and obtain updated attention weights for Eqn. 1.
4.3 Domain adaptation of LN affine parameters
In TTA, prior methods often conduct adaptation on pre-trained models with batch normalization (BN) layers[22] and most of them are built upon BN statistics adaptation[43, 52, 56]. However, recent works[32, 5, 24] found out that layer normalization (LN, [2]) can be a better adaptation choice for transformers. Specifically, for a layer with $d$-dimensional input $x=(x^{(1)},\cdots,x^{(d)})^{\intercal}\in\mathbb{R}^{d}$, the layer normalized output are
$y^{(k)}=\gamma^{(k)}\hat{x}^{(k)}+\beta^{(k)},\text{i.e., }y=\gamma\odot\hat{x}+\beta,$ (3)
where $\hat{x}^{(k)}=\frac{(x^{(k)}-\mu_{d})}{\sqrt{\sigma_{d}^{2}+\epsilon_{\sigma}}}$, $\mu_{d}=\frac{1}{d}\sum_{k=1}^{d}x^{(k)},\\ \sigma_{d}=\frac{1}{d}\sum_{k=1}^{d}(x^{(k)}-\mu_{d})^{2}$. Here, $\gamma^{(k)}$ and $\beta^{(k)}$ are learnable affine parameters (and we denote the set of them as $\lambda$), $\epsilon_{\sigma}$ is a small value to prevent numerical problems. Adaptation to new domains can be conducted by tuning $\gamma^{(k)}$ and $\beta^{(k)}$ in Eqn. 3.
4.4 Sharpness-aware domain adaptation
While TTA provides stability in LN models, it can lead to model collapse during affine parameter tuning (Eqn. 3). This collapse occurs when the model incorrectly assigns all input samples to a single class over the adaptation process[35]. To prevent this, we utilize sharpness-aware techniques [10, 35] to make the model less sensitive to large gradients and in test samples [34].
Upon obtaining the updated weights from the updated module selector $\mathcal{G^{*}}$ for each test sample in the target batch $D_{T}^{(t)}=\{x_{i}^{(t)}\}_{i=1}^{B}$ at time step $t$, we identify the $M$ source modules most relevant to the target batch based on the highest average weights assigned by the module selector. We compute the entropy of the pseudo-labels of current test batch as predicted by these models with pretrained module. The entropy of the prediction of $t$-th target batch from $j$-th source model is:
$\mathcal{L}_{j}^{(t)}=-\mathbf{E}_{\mathcal{D}_{T}^{(t)}}\sum_{c=1}^{K}\hat{y}_{jc}^{(t)}\log(\hat{y}_{jc}^{(t)})$ (4)
We seek to make the model insensitive to the large gradients by encouraging the model to converge to a flat area of the entropy loss surface, since a flat minima has good generalization and robustness to large gradients[10, 35]:
$\min_{\lambda}\mathcal{L}_{j}^{SA(t)}(\{x_{i}^{(t)}\}_{i=1}^{B};\mathbf{\lambda}),$ (5)
where $\text{where }\mathcal{L}_{j}^{SA(t)}\triangleq\max_{\|\mathbf{\epsilon}\|_{2}\leq\rho}\mathcal{L}_{j}^{(t)}(\{x_{i}^{(t)}\}_{i=1}^{B};\mathbf{\lambda}+\mathbf{\epsilon})$ (6)
The abbreviation "SA" stands for sharpness-aware. In this context, the inner optimization aims to discover a perturbation $\mathbf{\epsilon}$ of LayerNorm affine parameters $\mathbf{\lambda}$ within a Euclidean ball of radius $\rho$ that maximizes entropy. The degree of sharpness is measured by the maximum change in the Euclidean ball neighbourhood $N_{\rho}(\mathbf{\lambda})$. This bi-level problem incentivizes the optimization process to locate flat minima. Following SAM[10], we can approximately solve the inner optimization via first-order Taylor expansion
$\mathbf{\epsilon^{*}}(\mathbf{\lambda})$ $\triangleq\operatorname*{arg\,max}_{\|\mathbf{\epsilon}\|_{2}\leq\rho}\mathcal{L}_{j}^{(t)}(\{x_{i}^{(t)}\}_{i=1}^{B};\mathbf{\lambda}+\mathbf{\epsilon})$
$\approx\operatorname*{arg\,max}_{\|\mathbf{\epsilon}\|_{2}\leq\rho}\mathcal{L}_{j}^{(t)}(\{x_{i}^{(t)}\}_{i=1}^{B};\mathbf{\lambda})+\mathbf{\epsilon}^{\intercal}\nabla_{\lambda}\mathcal{L}_{j}^{(t)}(\{x_{i}^{(t)}\}_{i=1}^{B};\mathbf{\lambda})$
$=\operatorname*{arg\,max}_{\|\mathbf{\epsilon}\|_{2}\leq\rho}\mathbf{\epsilon}^{\intercal}\nabla_{\lambda}\mathcal{L}_{j}^{(t)}(\{x_{i}^{(t)}\}_{i=1}^{B};\mathbf{\lambda})$
Denote $\mathbf{v}=\nabla_{\lambda}\mathcal{L}_{j}^{(t)}(\{x_{i}^{(t)}\}_{i=1}^{B};\mathbf{\lambda})$. Hölder’s inequality implies that $\epsilon^{\mathbf{T}}\mathbf{v}\leq\|\epsilon\|_{p}\|\mathbf{v}\|_{q}\leq\rho\|\mathbf{v}\|_{q}$ ($11/p+1/q=$ 1). For $p=q=2$, the linear function achieves that bound $\mathbf{\epsilon^{*}}(\mathbf{\lambda})^{\intercal}\mathbf{v}=\rho\|\mathbf{v}\|_{2}$, and
$\mathbf{\epsilon^{*}}(\mathbf{\lambda})=\rho\text{sgn}(\mathbf{v})\frac{|\mathbf{v}|}{\|\mathbf{v}\|_{2}}$ (7)
By substituting $\mathbf{\epsilon^{*}}(\mathbf{\lambda})$(Eqn. 7) back into Eqn. 6 and differentiating both sides, the final gradient approximation is:
$\nabla_{\lambda}\mathcal{L}_{j}^{SA(t)}\approx\nabla_{\lambda}\mathcal{L}_{j}^{(t)}(\{x_{i}^{(t)}\}_{i=1}^{B};\mathbf{\lambda})|_{\lambda+\mathbf{\epsilon^{*}(\mathbf{\lambda})}}.$ (8)
By applying Eqn. 8 to update the LN affine parameters of $j$-th model with Eqn. 4, we can obtain a more reliable solution for entropy minimization. We update all the LayerNorm affine parameters of $j$-th source model by back propagating $\mathcal{L}_{j}^{(t)}$ once, as the approach in [52]. After updating $\gamma$s and $\beta$s of LN layers of $j$-th model, we denote the updated model as $\overline{f_{\theta_{j}}^{j(t)}}$ which will be used of the inference on $(t+1)$-th target test batch. This implies $f_{\theta_{j}}^{j(t+1)}=\overline{f_{\theta_{j}}^{j(t)}}$.
5 Experiment
Table 2: Results on Office-Home. We train the source models using 3 datasets to perform inference on the remaining dataset. The table clearly shows that PLUTO outperforms all of the baselines consistently
Source Method Ar Cl Pr Re Avg
Single Source-Best 62.5 47.7 74.3 69.4 63.4
Source-Worst 42.5 40.7 58.2 60.1 50.4
TENT-Best 65.5 51.2 76.8 71.4 66.2
TENT-Worst 46.6 45.7 56.3 64.0 53.2
Multi TENT-Ens 63.1 49.2 75.6 71.0 64.7
TENT-Ens-SA 64.1 50.8 77.1 74.5 66.6
PLUTO (Ours) 66.0 53.6 79.2 76.3 68.7
Table 3: Results on Digits. We train the source models using 4 digits datasets to perform inference on the remaining dataset. The table clearly shows that PLUTO outperforms all of the baselines consistently
Source Method MM MT UP SV SY Avg
Single Source-Best 53.5 98.4 84.5 83.0 94.5 82.7
Source-Worst 21.4 42.2 51.6 12.3 17.1 28.9
TENT-Best 56.1 98.4 84.9 87 95.2 84.3
TENT-Worst 17.6 54.2 59.6 11.4 15.5 31.6
Multi TENT-Ens 62.4 97.8 87.6 55.1 78.4 76.2
TENT-Ens-SA 62.8 98.8 88.7 71.2 86.7 81.6
PLUTO (Ours) 63.1 98.9 89.6 86.4 93.7 86.4
Table 4: Diminishing return. On ImageNet-C (severity level=2) dataset, we train 15 modules for 15 domains, along with the module selector $\mathcal{G}$. For the test-time adaptation performance evaluation, we pick 14 of them to adapt to the rest one. $M$ is the number of selected module, $N$ is the size of the set of all modules. We report the averaged accuracy over 15 target domains for different choice of $M$.
Description $M=1$ $M=3$ $M=5$ $M=8$ $M=11$ $M=14=N-1$
Unlabeled update to module selector $\mathcal{G}$, Adaptation with $M$ modules selected 56.9 59.4 63.7 64.1 64.3 64.4
No update to module selector $\mathcal{G}$, Zero-shot select $M$ source modules, No LN tuning adaptation 42.5 47.3 51.3 53.1 53.3 53.5
Table 5: Larger module store. On ImageNet-C dataset, there are 15 different corruptions with 5 different severity levels, leading to at most $5\times 15=75$ different domains. For the test-time adaptation performance evaluation, we pick one target domain and select $M=14$ source modules to adapt to it. We evaluated the module selection performance on ImageNet-C dataset. Note that, given more source domains for pretraining, for larger $N$, the module selector $\mathcal{G}$ is trained on more source domains, which leads to more complexity. We report the averaged accuracy over $N$ target domains for different choice of $N$.
Description $N=15$ $N=30$ $N=50$ $N=75$
Unlabeled update to module selector $\mathcal{G}$, Adaptation with $M=14$ modules selected 64.4 65.5 65.8 66.2
No update to module selector $\mathcal{G}$, Zero-shot select $M=14$ source modules, No LN tuning adaptation 53.5 55.4 56.5 57.2
Table 6: Few-shot performance of PLUTO. On ImageNet-C (severity level=2) dataset, we train 15 modules for 15 domains. For the test-time adaptation performance evaluation, we pick 14 of them to adapt to the rest one. We randomly pick $U<B$ samples in the test batch to update module selector, make module selection, and LN tuning adaptation. We report the averaged accuracy over 15 target domains for different choice of $U$. $\mathcal{G}$ is pretrained on 15 source domains in a supervised manner.
Description $U=B=128$ $U=64,B=128$ $U=32,B=128$ $U=16,B=128$ $U=4,B=128$ $U=0$, Zero-shot
Baseline Fully-finetune on $T_{target}$ 76.7 $N/A$ $N/A$ $N/A$ $N/A$ $N/A$
Supervised PET on $T_{target}$ 72.1 $N/A$ $N/A$ $N/A$ $N/A$ $N/A$
Randomly select $M=5$ modules, No adaptation $N/A$ $N/A$ $N/A$ $N/A$ $N/A$ 42.2
PLUTO Supervised update to module selector $\mathcal{G}$ on $T_{target}$, Adaptation with $N=14$ modules 68.2 67.5 67.1 65.3 63.1 $N/A$
Unlabeled update to module selector $\mathcal{G}$, Adaptation with $M=N=14$ modules selected 64.4 63.1 62.2 60.9 57.4 50.1
Unlabeled update to module selector $\mathcal{G}$, Adaptation with M=5 modules selected 63.7 62.2 61.2 59.4 56.3 47.3
Table 7: Comparing the issue of forgetting to the baselines on CIFAR-10, this table shows that PLUTO displays significantly slower forgetting rates over an extended cycle of repetitive test distributions in comparison to the baseline methods. It’s worth noting that the "online" methods, which rely on additional domain knowledge, serve as upper bounds for their respective continuous adaptation counterparts.
Time $t$
Round 1 3 5 All
Conditions Clean Snow Fog Frost Clean Snow Fog Frost Clean Snow Fog Frost Avg
TENT-Best-Online 89.9 87.7 87.6 90.4 89.9 87.7 87.6 90.4 89.9 87.7 87.6 90.4 88.9
TENT-Ens-Online 94.1 91.1 93.8 92.5 94.1 91.1 93.8 92.5 94.1 91.1 93.8 92.5 92.9
TENT-Ens-SA-Online 94.3 91.4 94 93.1 94.3 91.4 94 93.1 94.3 91.4 94 93.1 93.2
PLUTO-Online 95.1 93.3 94.4 94.4 95.1 93.3 94.4 94.4 95.1 93.3 94.4 94.4 94.3
TENT-Best 89.9 78.9 82.1 70.2 53.7 45.8 50.6 47 40.7 29.7 31.1 27.4 53.9
TENT-Ens 94.1 85.2 83.1 77.5 61.8 54.4 52.3 47.5 35.2 30.4 30.8 28.7 56.7
TENT-Ens-SA 94.3 88.2 83.7 80.6 66.7 57.5 59.7 62.8 57.4 40.6 30.8 43.1 63.7
PLUTO 95.1 92.5 93.9 92.2 89.1 92.6 89.5 88.6 87.3 87.8 86.9 87.9 90.3
In the main paper, we employ parameter-efficient fine-tuning (PET) in the form of VPT [23, 13] as our module to demonstrate the effectiveness of our approach. We will supplement our method’s application to other PET modules in the supplementary material to illustrate the generality of our approach. The effectiveness of our approach is evaluated on both stationary and continuous test time adaptation benchmarks.During the testing phase, we plan to evaluate our approach under two scenarios: (1) Test batches are sampled from a stationary distribution (2) Test batches are sampled from a evolving distribution.
Datasets. We demonstrated the efficacy of our approach in handling stationary distributions by assessing its performance on the following domain adaptation benchmarks.
  • Office-Home [50] The Office-Home dataset consists of four domains: Art (Ar), Clipart (Cl), Product (Pr), and Real World (Re), each containing 65 classes. For evaluating our method in the test-time adaptation scenario, we selected three of these domains as source domains and used the remaining domain as the test data. We conducted online inference on the test data in batches using the source domains, while the test data had its labels removed.
  • ImageNet-C. [17] ImageNet-C dataset consists of 15 diverse corruption types applied to validation images of ImageNet[6]. It consists algorithmically generated corruptions from noise, blur, weather, and digital categories with 5 severity levels. We regard different corruptions with different severity level as different domains. Thus we will have $5\times 15=75$ domains at most.
  • Digits-Five. [36] consists of five-digit datasets: MNIST (mt), MNIST-M (mm), USPS (up), SVHN (sv), Synthetic Digits (sy). There are 10 classes corresponding to digits ranging from 0 to 9. We employed four of the domains as sources and reserved the remaining domain for testing.
In the context of continuous test-time adaptation, we employ the following benchmarks.
  • CIFAR-10C. [17] The CIFAR-10[26] dataset is an image classification dataset consisting of 50,000 training images and 10,000 test images. Based on CIFAR-10, CIFAR-10C [17], introduced 15 different types of noise at different severity levels (1 $\rightarrow$ 5) to the 10,000 test images. This dataset has since been utilized in various studies related to continuous test-time adaptation.
5.1 Baseline Methods
Our main baseline is TENT[52], which is designed for single-source test-time adaptation. We also explore some variants of TENT. To assess adaptation performance, we adopt a setup similar to that in [1]. We apply each source model to the test data independently, resulting in TENT-Best and TENT-Worst, representing the best and worst adaptation performances among these models, respectively. The best/worst results of directly applying source models on target test data without adaptation are noted as Source-Best and Source-Worst. Additionally, we calculate the average pre-softmax prediction (ensemble) over all sources and refer to it as TENT-Ens. To make fair comparison, we also have sharpness-aware (SA) version for the above baselines.
For adaptation on stationary distributions, as our method is based on the ensemble of pre-trained ViT models with pre-trained modules on unlabeled data, we will compare our approach with the following baseline methods: (1) Full model fine-tuning of a pretrained ViT model. (2) Freezing most parameters of a pretrained ViT model and supervising training a module $\theta_{T}$ on the target domain. This comparison will help evaluate the effectiveness of our method against these different fine-tuning/training approaches.
For adaptation on continuous test-time adaptation, to emphasize the issue of forgetting, we contrast our approach with the online variants of the mentioned methods. In the online scenario, we have access to an oracle that signals domain changes, presenting an ideal situation. Here, the source models are reset whenever the oracle identifies a new domain. However, it just serves as an upper-bound solely for continuous adaptation scenarios where domain knowledge is accessible, which is not a practical real-world scenario.
5.2 Implementation details
We use ViT-Base-16[8] model pretrained on JFT-300M dataset for all our experiments. For all the experiments we use a target batch size of $B=128$, as used by TENT [52]. For more details on implementation and experimental setting see supplementary.
5.3 Object recognition and digit classification
We report the results of object recognition on Office-home [50] and Digits-Five [36] dataset in Table. 23. We compute the accuracy for each incoming test batch and subsequently present the results by averaging the accuracy values across all the batches. For Office-home dataset, the number of source modules is $N=3$, and we use all $M=N$ (in Alg. 1) to do the inference on the target domain. Similar setup for Digits-Five dataset, except $M=N=4$ here.
5.4 Module selection
In the practical application scenarios of our method, there may be a considerable number of candidate pretrained modules corresponding to various source domains. However, not all of these source domains have a significant impact on the adaptation to the target domain. Therefore, as illustrated in Algorithm  1, out of all the $N$ pretrained source modules, we only utilize $M(<N)$ modules for domain adaptation. The core idea of our approach is: (1) to identify the most relevant source domains/source modules to the target domain and (2) to assign greater weights to these modules in the model ensemble to provide better-weighted pre-softmax logits. We will demonstrate that we have achieved these two points through experiments from two aspects:
(1) Diminishing return. We select a growing number of modules from a fixed set: We believe that by measuring the average weight allocated to each source by the module selector on a target test batch, we can reflect the general importance of these source domains to this batch’s relationships. Therefore, by selecting the top $M(M<N)$ source modules corresponding to the highest average weights and using only these M sources for adaptation.
(2) Larger module store Adding more pretrained modules (leading to larger $N$) to the current module set while keeping the number of modules involved in adaptation the same. PLUTO can still identify the most important modules while the size of the module store increases.
5.5 Few-shot adaptation
As shown in Alg. 1, our approach involves training an atTENTion-style module selector responsible for generating weights to ensemble the outputs of various source models. This process is designed to tailor predictions for each target sample. By doing so, PLUTO can effectively account for sample-specific preferences in ensembling different source models, particularly when provided with limited labeled or unlabeled target data, as shown in Table. 6
We showcase the few-shot performance via demostrating results on ImageNet-C. On ImageNet-C dataset, we train 15 modules for 15 corruption domains (with severity level=2). Then for $t$-th target batch $D_{T}^{(t)}=\{x^{(t)}_{i}\}_{i=1}^{B}$ Use we pick different $U(<B)$ samples to update module selector $\mathcal{G}$ and make selection in Alg. 1.
5.6 Forgetting issue
We extensively experiment with this dataset to assess our model’s performance in the face of dynamic test distributions. We introduce three types of corruption noise from CIFAR-10C: Snow, Fog, and Frost, each with a severity level of 5. These noises are added to the original CIFAR-10 training set, leading to the training of three source modules, each dedicated to one of these noise types. Additionally, we train a module on the clean training set of CIFAR-10. In other words, we trained $N=4$ modules on noisy CIFAR-10C dataset. Evaluation of forgetting conducted on other datasets will be included in the supplementary material.
During testing, we evaluate these models sequentially on test data corresponding to each noise type. We progressively increase the severity level until reaching the maximum, followed by a gradual decrease in severity level in reverse order. The sequence structure is as follows: (Snow) severity: $\cdots\rightarrow 1\rightarrow 2\rightarrow 3\rightarrow 4\rightarrow 5\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow\cdots$ (Fog) severity:$\cdots\rightarrow 1\rightarrow 2\rightarrow 3\rightarrow 4\rightarrow 5\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow\cdots$, and so forth.
We repeat this entire cycle five times to observe how each model’s performance deteriorates over time due to forgetting. In each of the five cycles, the test data remain identical, allowing for a fair comparison with previous cycles. We present the results of this experiment in Table.7, distinguishing between the online and continual versions with a horizontal line. The online version represents the best-case scenario that relies on additional oracle domain information.
6 Conclusion
We present PLUTO, a Test Time Adaptation (TTA) algorithm for transformers, combining multiple source parameter-efficient tuning (PET) modules during test time. PLUTO achieves test accuracy exceeding the best individual source model and prevents catastrophic forgetting. Experiments reveal two key findings: (i) PLUTO outperforms TENT-Best, TENT-Ens, and TENT-Ens-SA for various test distributions and achieves good few-shot performance. (ii) Over repeated cycles, baseline methods degrade significantly, while PLUTO maintains high performance with slower forgetting.